How many pounds of hamburger that costs $1.60 per pound must be mixed with 70 pounds of hamburger that costs $2.10 per pound to make a mixture that costs $1.70 per pound.
Answer by stanbon(48558)

How many pounds of hamburger that costs $1.60 per pound must be mixed with 70 pounds of hamburger that costs $2.10 per pound to make a mixture that costs $1.70 per pound.
------------------------------
Equation:
value + value = value
1.6x + 2.1*70 = 1.7(x+70)
Multiply thru by 10 to get:
16x + 21*70 = 17x + 17*70
x = 4*70
x = 280 lbs. (amt. of $1.60 hamburger in needed in the mixture)
--------
Cheers,
Stan H.